- Status Report
- Jan 29, 2023
NOAA GOES SATOPS Morning Report 31 Jan 2001
SATOPS MORNING REPORT – January 31, 2001
No change in the status of the POES spacecraft.
NOAA-16 rev 1853 / W at 1928Z on January 30: L5B short due to
intermittent drop-outs; 33.4% recovered. Further recovery attempts
not possible due to spacecraft recorder management constraints.
NOAA-16 was returned to the secure (encrypted) mode of commanding during
rev 1849 / F.
To support troubleshooting of the NOAA-16 CPU/CIU timing problem, a
test sequence will be commanded to NOAA-16 at 12:38Z (rev 1877 / F, 76
deg elevation) on February 1.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 1731Z on January 30, a GOES-8 E/W station keeping maneuver was successfully
executed. The maneuver duration was 6.1 seconds and was completely
nominal. Two attempts at transition mode (from thruster to wheel
control mode) were executed in an attempt to minimize initial transition
At 1926Z on January 30, an imager frame break was observed on GOES-8.
This frame break was attributed to loss of signal synchronization.
The GOES-11 storage mode attitude has deviated from the proper spin
axis to the extent that a ZAP re-spin maneuver is planned for February
1 (approximately 11:00pm to 2:00am EST tonight), to reset the spin axis
on the Sun.
GOES operations were nominal over the past 24 hours.
Questions or comments to [email protected]